OpenStudy (anonymous):

Find a trinomial for the area of the rectangular rug shown whose sides are x+5 feet(height) and 2x-3 feet(length)

OpenStudy (anonymous):

= (x+5)(2x-3) = 2x^2 + 7x = 15

OpenStudy (anonymous):

do you need any more explanation uriel?

OpenStudy (anonymous):

I'm little lost on the answer since the answer on the book is(x2(square)+4x+4) in2(square) but I only get the same answer: (2x2(square)+7-3. Still thank you for your help.

OpenStudy (anonymous):

i made a typo - it should be 2x^2 + 7x - 15

OpenStudy (anonymous):

oh I see now it's my fault. Sorry Jimmy. I read the answer wrong. Still thank you for your help.